OpenStudy (anonymous):

dissolve 3g of sodium nitrate in water to make 250 ml of solution. The density of this solution is equal to 1.12g/mL. Determine the molar, normal, molal and percentage mass concentration of the solution

OpenStudy (anonymous):

n= 3g/85g= 0.035mol V=250ml/1000= 0.25 d= 1.12g/ml molar= n/V = 0.035/0.25 molar= 0.14M